# references.bib

### v5

@@ 4.3 Symbolic policies generalise to varying environments
Besides partial observations, realistic settings also require policies to generalise to varying and changing environments. To validate that generalising symbolic policies can be evolved, we test policies in the SHO with parameters that vary between trials (Fig. 4e). In this experiment, the upper bound fitness is again obtained with the LQG controller that is computed independently for each parameter setting (black curves). The NDE converges close to the upper bound fitness and is consistent across different runs (green curve and dots).
Static policies are evolved to validate their generalisability (blue curve). GP-S does not converge near the upper bound fitness and the NDE. The validation fitness of the best static policies shows that a ceiling has been reached, as every policy is located close to the average (blue dots). GP-D converges to a similar fitness as GP-S (orange curve). However, the validation fitness of the dynamic symbolic policies shows an interesting pattern (orange dots). Many dynamic symbolic policies obtain a validation fitness that is better than any static policy, and some even compare with NDE policies. However, some dynamic symbolic policies obtain a validation fitness worse than any static policy.
The inconsistent evolution of dynamic symbolic policies is explained by the large search space. When dynamic policies are evolved successfully, they perform better than static policies. The memory allows dynamic policies to discover useful properties that increase generalisability. In our results, the majority of evolutionary runs resulted in dynamic policies that exceed the performance of static policies. Therefore the addition of the latent state in symbolic policies proves to be more beneficial on average. Our approach does not consistently produce policies that perform comparably as the NDE policies, though this comes at the expense of obtaining white-box solutions.
@@ 4.4 Genetic programming evolves higher-dimensional policies
As the search space of symbolic policies is large, especially for the dynamic symbolic policies, it is important to validate whether high-performing policies are found efficiently when the control dimension increases. In this experiment, the policies can apply control forces to both links to accomplish the acrobot swing-up (Fig. 4f). The NDE converges to a slightly higher fitness during evolution than in the other experiments on the acrobot, because more force can be applied to solve the swing up (green curve). The best NDE policies obtain consistent validation fitness (green dots).
Our method is first tested to produce static symbolic policies with two control outputs (blue curve). The search space of static policies is still small compared to the dynamic policies, as is observed from the higher initial average best fitness compared to the other methods. The average best fitness of GP-S does not improve much during evolution, and converges at a worse fitness than the NDE. The validation fitness shows that most static policies perform similarly, but one evolved policy performs better than the other static policies (blue dots). GP-D converges at a slightly better fitness than GP-S, but worse than NDE (orange curve). The validation fitness obtained by dynamic symbolic policies mostly overlaps with that of the static symbolic policies and NDE policies, although a few policies are worse than any policy obtained by the other methods (orange dots). This experiment demonstrates that GP still evolves higher-dimensional control policies that accomplish the acrobot swing up, even though the size of the search space was increased even further.
@@ 4.5 Generalising symbolic policies are discovered for an industrial application
In the final experiment, we apply our method to the continuous stirred tank reactor (CSTR) as an industrial problem with non-linear dynamics. The temperature in the reactor has to be stabilised at a setpoint temperature by controlling the flow of coolant into the cooling jacket (Fig. 5a). To increase the complexity of the problem, the environment setting is varied between trials and the policies receive noisy partial state observations.
The NDE has not converged after the dedicated number of generations, while being stuck at a poor fitness for the first half of the evolution (green curve). A potential reason for the limited convergence of NDEs could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently. The validation fitness of the NDE policies is on average lower than the symbolic policies, and there are a few bad-performing policies (green dots).
GP-S converges to a higher fitness than NDE (blue curve), and the validation fitness shows high consistency among the policies (blue dots). GP-D obtains a better fitness than GP-S during evolution (orange curve). GP-D discovered several policies that outperform GP-S and NDE in terms of validation fitness (orange dots). However, the high-performing dynamic symbolic policies are not consistently found, as some policies obtain a validation fitness significantly worse than any static policy.
@@
Figure 5: Dynamic symbolic policies generalise to industrial control settings. (a) Evolution results of the experiment on the continuous stirred tank reactor (CSTR) with partial state observability and varying parameters. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies and neural differential equation (NDE). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (b) Simulated trajectory of the CSTR controlled with a dynamic symbolic policy. The top row shows the temperature in the reactor and cooling jacket and the controlled coolant flowrate into the cooling jacket with varying setpoint temperatures. The bottom row shows the latent variables of the policy. The policy used in this simulation is described in Table 2 (CSTR, Dynamic). (c) The temperature of the reactant that flows into the main reactor increases throughout the trajectory. (d) The observations of the temperature are made more noisy. (e) The dynamics of the temperature in the main reactor are influenced by an external sinusoidal wave, simulating the change of temperature outside of the reactor during the day and night.
@@ 4.6 The latent variables in dynamic symbolic policies are transparent
Across the experiments, GP was able to find successful dynamic symbolic policies. The best validation fitness obtained with each method is presented in Table 1. GP-D is generally competitive with black-box NDEs, and outperforms the NDE in 3 out of 7 experiments. The best validation fitness of GP-D is higher than GP-S for every experiment. Table 1 also presents the complexity of the best policies in terms of the number of mathematical operations, variables and constants. GP-S always obtains policies with lower complexity than GP-D, which is partially accounted for by the fact that GP-D consists of three trees. The dynamic symbolic policies have an increased complexity compared to the static policies, but this complexity allows the policies to obtain higher performance in challenging experiments. Compared to the NDE, the complexity of GP-D is still significantly lower, while demonstrating competitive performance and more interpretability. The best static and dynamic symbolic policies are displayed in Table 2. Visualising the dynamic symbolic policies reveals that the latent variables exhibits transparent properties that allow us to investigate how the memory is used to solve control tasks.
@@
For this purpose, the best dynamic symbolic policy discovered with GP for stabilising the CSTR is analysed for its control and transparency (Table 2 CSTR, Dynamic). We interpret the state equation of variable $a_{1}$ as a scaling factor that converges to 2.92 during the trajectory. The state equation of variable $a_{2}$ computes the error between the temperature and setpoint, and adds a self-connection to improve robustness to noise. The error estimate is scaled by the sum of the control, $a_{2}$ and a constant. The readout scales $a_{2}$ with the squared of $a_{1}$, and adds a multiplication of $a_{1}$ with the logarithm of the setpoint temperature. The control therefore reacts to differences between the setpoint and current temperature, scaled increasingly during the trajectory.
This policy is subjected to a set of challenging experiments on the CSTR with varying parameter settings that were not included during evolution to demonstrate its generalisability and the flexibility of the latent state. During evolution, the setpoint temperature was fixed within a trajectory, but in Fig. 5b, the setpoint changes throughout the trajectory to test whether the policy can adjust its control. At the start of the trajectory, the temperature in the main reactor is quickly stabilised around the setpoint temperature without requiring much control (Fig. 5b, top row). However, when the setpoint temperature is changed, the policy accordingly adjusts its control signal to decrease the temperature to the setpoint. Similarly, the policy can successfully stabilise the temperature at each setpoint temperature. Latent variable $a_{1}$ converges to 2.92 as expected (Fig. 5b, bottom row, red curve), but $a_{2}$ shows large responses when the setpoint temperature changes (Fig. 5c, bottom row, blue curve), which is indirectly reflected in the control signal.
In industrial settings, not only the target may vary, but the environment itself can be volatile. To demonstrate the robustness of the dynamic symbolic policy, the temperature of the reactant that flows into the main reactor increases throughout the trajectory, which requires the policy to continuously adapt to satisfy the objective (Fig. 5c). The temperature in the main reactor increases constantly, but the policy increases its control signal, which consequently decreases the temperature in the cooling jacket and keeps the temperature in the reactor close to the setpoint (Fig. 5c, top row). Latent variable $a_{2}$ also increases over time, showing it reacts to the changing environment (Fig. 5c, bottom row). When the setpoint temperature is altered halfway, the latent state adjusts, and the control is temporarily decreased, after which is steadily increases again to match the changing feed temperature.
The previous two experiments tested for robustness to changing targets or environment settings. Another realistic scenario is that sensors deteriorate, and the observations may become less accurate. In Fig. 5d, the noise applied to the observations of the temperature in the reactor and coolant is increased by a factor of 4 compared to the setting during evolution. The policy still manages to stabilise the temperature around the setpoint, even after a change of the setpoint temperature (Fig. 5d, top row). The latent variable $a_{2}$ behaves more noisy (Fig. 5d, bottom row), which is also reflected in the control signal.
Finally, an external sinusoidal influence is added to the dynamics of the temperature in the reactor (Fig. 5e). The sinusoid resembles the oscillating temperature outside of the reactor during day and night. Although the policy stabilises the temperature around the setpoint, the oscillations in the temperature are still visible (Fig. 5e, top row). The control signal and the temperature in the cooling jacket show larger oscillations, which means the policy has to heavily adjust to compensate for the external influence. The latent variable $a_{2}$ also exhibits oscillatory behaviour, with a different pattern occurring after the setpoint has been changed (Fig. 5e, bottom row).
@@
These experiments demonstrate that the dynamic symbolic policies can generalise to challenging environment settings that were not used during evolution. Additionally, the equations of the dynamic symbolic policy are interpretable, and the latent variables offer transparency in the use of memory. The latent variables vary in their level of expressiveness, but overall analysing the latent variables aids with understanding the functionality of a policy.
@@ 5 Discussion
In this paper, we discovered interpretable and high-performing white-box control systems to solve control problems. The conventional approach is to find static symbolic policies, but these are difficult to extend to real-world challenges like partial observability and volatility. To improve the robustness of control systems, the policies have to be extended with memory. Before, memory was added to symbolic policies by providing time-lagged observations or allowing the policy to explicitly store and retrieve data. Here, we extend the policy with latent variables that are governed by differential equations. The latent variables continuously integrate observations in their dynamics, which allows for a more scalable and general approach compared to using time-lagged observations. By evolving symbolic state equations of the latent variables with genetic programming (GP), the resulting policies consist of interpretable equations and the memory is transparent.
The dynamic policies successfully solved control tasks with noisy and partial state observations, where the static policies obtain limited results. Furthermore, dynamic symbolic policies generalise better to tasks with varying environment settings than static policies. These results show the benefits of the memory in the dynamic symbolic policies, and these policies can still be evolved efficiently with GP in terms of generations and population size. The symbolic policies were compared with a neural differential equation optimised with covariance-matrix adaptation evolution strategies as a proven black-box baseline [salimans2017evolution]. The results demonstrate that dynamic symbolic policies are evolved that generally obtain comparable performance as the black-box baseline, although competitive symbolic policies are not consistently found in every evolutionary run.
@@
The major limitation of evolving dynamic symbolic policies with GP, is the huge search space that has to be traversed, which hinders consistent evolution of high-performing policies. The results showed that GP converges faster and more consistently when static policies are evolved under standard environment circumstances. The search space will grow even larger when the control problem contains higher-dimensional observation and control spaces. Our method demonstrated to be scalable when the environment required two control outputs (Fig. 4f), but this may not hold for more complex problems. Better policies can be evolved with GP when a higher population size and number of generations are selected, which also increases the computation requirements. However, previous research has focused extensively on improving the optimisation efficiency of GP, by integrating features such as optimisation of constants [topchy2001faster], semantic reproduction [vanneschi2014survey], population diversity control [burke2004diversity] and self-adaptation of hyperparameters [angeline1996two].
Besides introducing the concept of symbolic policies with dynamic latent memory, this research also resulted in a general GP framework that is not limited to optimisation of dynamic symbolic policies. First of all, other structures of symbolic policies can be configured, with the possibility to explicitly define different types of trees with unique inputs and functionality in the policy. Furthermore, the framework could also be used problems outside control, for example symbolic regression [bongard2007automated], discovering learning rules [jordan2021evolving] or other directions within scientific discovery [wang2023scientific, vinuesa2024opportunities, rackauckas2020universal].

### v6

@@ 4.3 Symbolic policies generalise to varying environments
Besides partial observations, realistic settings also require policies to generalise to varying and changing environments. To validate that generalising symbolic policies can be evolved, we tested policies in the SHO with parameters that vary between trials (Fig. 3d). In this experiment, the upper bound fitness was again obtained with the LQG controller that is computed independently for each parameter setting (horizontal black lines). The NDE converged close to the upper bound fitness and is consistent across different runs (green curve and dots).
Static policies were evolved to validate their generalisability (blue curve). GP-S did not converge near the upper bound fitness and the NDE. The validation fitness of the best static policies shows that a ceiling has been reached, as every policy is located close to the average (blue dots). GP-D converged to a similar fitness as GP-S (orange curve). However, the validation fitness of the dynamic symbolic policies shows an interesting pattern (orange dots). Many dynamic symbolic policies obtain a validation fitness that is better than any static policy, and some even compare with NDE policies. However, some dynamic symbolic policies obtain a validation fitness worse than any static policy.
Appendix Fig. 7a shows that most policies perform similarly on the test data as on their respective training data sets. However, some policies fail to generalise, which is mostly the case for policies that include one or multiple non-linear operations. As the harmonic oscillator is governed by linear dynamics, the presence of non-linear operations in a policy could relate to overfitting on the training data and consequently harms the ability to generalise. The worst policy in terms of validation fitness is however extremely simple, as it does not make use of the latent memory.
The inconsistent evolution of dynamic symbolic policies is explained by the large search space and overfitting. When dynamic policies are evolved successfully, they perform better than static policies. The memory allows dynamic policies to discover useful properties that increase generalisability. In our results, the majority of evolutionary runs resulted in dynamic policies that exceed the performance of static policies. Therefore the addition of the latent state in symbolic policies proves to be more beneficial on average. Our approach does not consistently produce policies that perform comparably as the NDE policies, though this comes at the expense of obtaining white-box solutions.
@@ 4.4 Genetic programming evolves higher-dimensional policies
As the search space of symbolic policies is large, especially for the dynamic symbolic policies, it is important to validate whether high-performing policies are found efficiently when the control dimension increases. In this experiment, the policies can apply control forces to both links to accomplish the acrobot swing-up (Fig. 5d). The NDE converged to a slightly higher fitness during evolution than in the other experiments on the acrobot, because more force could be applied to solve the swing up (green curve). The best NDE policies obtained consistent validation fitness (green dots).
Our method was first tested to produce static symbolic policies with two outputs (blue curve). The search space of static policies is still small compared to the dynamic policies, as is observed from the higher initial average best fitness compared to the other methods. The average best fitness of GP-S did not improve much during evolution, and converged at a worse fitness than the NDE. The validation fitness shows that most static policies perform similarly, but one evolved policy performs better than the other static policies (blue dots). GP-D converged at a slightly better fitness than GP-S, but worse than NDE (orange curve). The validation fitness obtained by dynamic symbolic policies mostly overlaps with that of the static symbolic policies and NDE policies, although a few policies are worse than any policy obtained by the other methods (orange dots). This experiment demonstrates that GP still evolves higher-dimensional control policies that accomplish the acrobot swing up, even though the size of the search space was increased even further.
@@ 4.5 Generalising symbolic policies are discovered for an industrial application
In the final experiment, we applied our method to the continuous stirred tank reactor (CSTR) as an industrial problem with non-linear dynamics. The temperature in the reactor has to be stabilised at a setpoint temperature by controlling the flow of coolant into the cooling jacket (Fig. 6b). To increase the complexity of the problem, the environment setting is varied between trials and the policies receive noisy partial state observations.
The NDE did not converge after the dedicated number of generations, while being stuck at a poor fitness for the first half of the evolution (green curve). A potential reason for the limited convergence of NDEs could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently. The validation fitness of the NDE policies is on average lower than the symbolic policies, and there are a few bad-performing policies (green dots).
GP-S converged to a higher fitness than NDE (blue curve), and the validation fitness shows high consistency among the policies (blue dots). GP-D obtained a better fitness than GP-S during evolution (orange curve). GP-D discovered several policies that outperform GP-S and NDE in terms of validation fitness (orange dots). However, the high-performing dynamic symbolic policies were not consistently found, as some policies obtained a validation fitness significantly worse than any static policy.
Appendix Fig. 7b presents the fitness of the best dynamic symbolic policies on the training and testing data in the continuous stirred tank reactor experiment. In this experiment, the relation between the fitness obtained during evolution and the validation fitness shows a similar negative trend for all policies. The environment has a high degree of freedom, as the initial conditions, the setpoint temperature and nine parameters are sampled randomly in every trajectory. The poor generalisation of the policies could be explained by the fact that the training data is unable to fully represent the space of possible experiment settings, and therefore the policies are tested in unfamiliar settings during validation.
@@
Figure 6: Dynamic symbolic policies generalise to industrial control settings. (a) The continuous stirred tank reactor (CSTR) is described by the concentration of reactant $c$, the temperature $T_{r}$ in the main reactor and the temperature $T_{c}$ in the cooling jacket. The temperature in the reactor should be stabilised at a set point temperature $T_{r}^{*}$. The flow of coolant into the cooling jacket is controlled by $u$. (b) Evolution results of the experiment on the CSTR with partial state observability and varying parameters. The left plot shows the best fitness at every generation during evolution, averaged over 20 independent runs. The methods included are genetic programming for static (GP-S) and dynamic (GP-D) policies and neural differential equation (NDE). The right plot presents the validation fitness obtained on unseen conditions with the best policy evolved in every evolutionary run. (c) Simulated trajectory of the CSTR controlled with a dynamic symbolic policy. The top row shows the temperature in the reactor and cooling jacket and the controlled coolant flowrate into the cooling jacket with varying setpoint temperatures. The bottom row shows the latent variables of the policy. The policy used in this simulation is described in Table 2 (CSTR, Dynamic). (d) The temperature of the reactant that flows into the main reactor increases throughout the trajectory. (e) The observations of the temperature are made more noisy. (f) The dynamics of the temperature in the main reactor are influenced by an external sinusoidal wave, simulating the change of temperature outside of the reactor during the day and night.
@@ 4.6 The latent variables in dynamic symbolic policies are transparent
Across the experiments, GP was able to find successful dynamic symbolic policies. The best validation fitness obtained with each method is presented in Table 1, together with the complexity of the best policy. GP-D is generally competitive with black-box NDEs, and outperforms the NDE in 3 out of 7 experiments. The best validation fitness of GP-D is higher than GP-S for every experiment. Table 1 also presents the complexity of the best policies in terms of the number of mathematical operations, variables and constants. GP-S always obtains policies with lower complexity than GP-D, which is partially accounted for by the fact that GP-D consists of three trees. The dynamic symbolic policies have an increased complexity compared to the static policies, but this complexity allows the policies to perform better in challenging experiments. Compared to the NDE, the complexity of GP-D is still significantly lower, while demonstrating competitive performance and more interpretability.
To understand the workings of a dynamic symbolic policy, we can inspect the mathematical equations and visualise the latent memory of such a policy. The best static and dynamic symbolic policies are presented in Table 2. Generally, the dynamic symbolic policies consist of few variables and the equations are easily readable. The policies include sub-expressions for computing the distance to the target and state estimation from noisy observations. Nonetheless, it does not immediately become trivial to fully understand a dynamic symbolic policy purely from the equations. To this end, the latent memory can be visualised to inspect the policy’s behaviour in a trajectory.
@@
As a demonstration, the best dynamic symbolic policy discovered with GP for stabilising the CSTR is analysed for its control and transparency (Table 2 CSTR, Dynamic). We interpret the state equation of variable $a_{1}$ as a scaling factor that converges to 2.92 during the trajectory. The state equation of variable $a_{2}$ computes the error between the temperature and setpoint, and adds a self-connection to improve robustness to noise. The error estimate is scaled by the sum of the control, $a_{2}$ and a constant. The readout scales $a_{2}$ with the squared of $a_{1}$, and adds a multiplication of $a_{1}$ with the logarithm of the setpoint temperature. The control therefore reacts to differences between the setpoint and current temperature, scaled increasingly during the trajectory.
This policy is subjected to a set of challenging experiments on the CSTR with varying parameter settings that were not included during evolution to demonstrate its generalisability and the flexibility of the latent state. During evolution, the setpoint temperature and parameters of the environment were fixed within a trajectory, but in these example trajectories they were varied. The policy adjusted its control signal accordingly when the target changed multiple times throughout a trajectory (Fig. 6c, green curve), when the temperature feed into the reactor changed over time (Fig. 6d, green curve), with higher observation noise than it was evolved for (Fig. 6e, green curve) and with an external sinusoidal influence that resembles the change of outside temperature during day and night (Fig. 6f, green curve). In each trajectory, the policy shows a different control strategy, displaying its robustness and flexibility.
Besides robustness, these experiments also show high transparency in the control strategy to the different circumstances. The applied control signal is a mapping of the latent state that integrates the observations. As expected, the latent variable $a_{1}$ converges to 2.92 in the trajectories (Fig. 6c, red curve). Latent variable $a_{2}$ is more expressive and shows adaptations in each experiment. In Fig. 6c (blue curve), the latent state stabilised over time, but changed once a new setpoint temperature is required. Similarly, the latent state increased over time as the feed temperature rises (Fig. 6d, blue curve). With higher observation noise, the latent state shows more stochasticity (Fig. 6e, blue curve), which is also portrayed in the control signal. Finally, the latent state oscillates to compensate for the external sinusoidal influence, with a different pattern appearing after the setpoint temperature changes (Fig. 6f, blue curve).
@@
These experiments demonstrated that the dynamic symbolic policies can generalise to challenging environment settings that were not used during evolution. Additionally, the equations of the dynamic symbolic policy are interpretable, and the latent variables offer transparency in the use of memory. The latent variables vary in their level of expressiveness, but overall analysing the latent variables aids with understanding the functionality of a policy.
@@ 5 Discussion
In this paper, we presented a method to discover interpretable and high-performing white-box control systems to solve control problems. The conventional approach is to find static symbolic policies, but these are difficult to extend to real-world challenges like partial observability and volatility. To improve the robustness of control systems, the policies have to be extended with memory. Before, memory was added to symbolic policies by providing time-lagged observations [hein2018interpretable, hein2021trustworthy] or allowing the policy to explicitly store and retrieve data [koza1999genetic]. Here, we extend the policy with latent variables that are governed by differential equations. The latent variables continuously integrate observations in their dynamics, which allows for a more scalable and general approach compared to using time-lagged observations [khan2017memory]. By evolving symbolic state equations of the latent variables with genetic programming (GP), the resulting policies consist of interpretable equations and the memory is transparent.
The dynamic policies successfully solved control tasks with noisy and partial state observations, where the static policies obtained limited results. Furthermore, dynamic symbolic policies generalised better to tasks with varying environment settings than static policies. These results showed the benefits of the memory in the dynamic symbolic policies, and these policies can still be evolved efficiently with GP in terms of generations and population size. The symbolic policies were compared with a neural differential equation optimised with covariance-matrix adaptation evolution strategies as a proven black-box baseline [salimans2017evolution]. The results demonstrated that dynamic symbolic policies were evolved that generally obtain comparable performance as the black-box baseline, although competitive symbolic policies were not consistently found in every evolutionary run.
@@
A limitation of evolving dynamic symbolic policies with GP is that high-performing policies are not consistently discovered. Appendix Fig. 7 shows that some discovered policies displayed poor generalisation to unseen cases. In the harmonic oscillator, the limited generalisation could be partially explained by the occurrence of overfitting in some runs, related to evolving non-linear functions. In the continuous stirred tank reactor, all policies showed a decrease of performance on the validation trajectories, which is the result of the high number of free parameters in the environment. Overfitting is however a general phenomenon and not a inherent limitation of the dynamic symbolic policies or GP itself, and it shows that the introduced method is prone to overfitting when the data does not contain enough variation.
Another factor that hinders consistent evolution of high-performing policies, is the large search space that has to be traversed for the dynamic symbolic policies. The results showed that GP converges faster and more consistently when static policies were evolved under standard environment circumstances. The search space will grow even larger when the control problem contains higher-dimensional observation and control spaces. Our method demonstrates to be scalable when the environment required two control inputs (Fig. 5d), but this may not hold for more complex problems. Better policies can be evolved with GP when a higher population size and number of generations are selected, which also increases the computation requirements. However, previous research has focused extensively on improving the optimisation efficiency of GP, by integrating features such as optimisation of constants [topchy2001faster], semantic reproduction [vanneschi2014survey], population diversity control [burke2004diversity] and self-adaptation of hyperparameters [angeline1996two].
Besides introducing the concept of symbolic policies with dynamic latent memory, this research also resulted in a general GP framework [de2025kozax] that is not limited to optimisation of dynamic symbolic policies. First of all, other structures of symbolic policies can be configured, with the possibility to explicitly define different types of trees with unique inputs and functionality in the policy. Furthermore, the framework could also be used problems outside control, for example symbolic regression [bongard2007automated], discovering learning rules [jordan2021evolving] or other directions within scientific discovery [wang2023scientific, vinuesa2024opportunities, rackauckas2020universal].